OpenStudy (anonymous):

Help please...Solve the following problem by writing a system of equations: A 25% alcohol solution is to be mixed with a 45% alcohol solution to obtain 8 liters of a 35% alcohol solution. How many liters of each solution should be used?

OpenStudy (anonymous):

65%x=35%(8) 65/100x=2.8 x=4.3 liters

OpenStudy (anonymous):

Wait a min I think im wrong (25%+45%)x=35%(8) .7x=2.8 X=4 liters At first I added wrong

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

lets call \(x\) the amount of 45% solution, so it will contain \(.45x\) alcohol then the amount of 25% solution must be \(8-x\) because the total is 8 it will contain \(.25(x-8)\) alcohol the total will therefore be \[.45x+.25(8-x)\] and you want that to be 35% of 8 which is 2.8 so set \[.45x+.25(8-x)=2.8\] and solve for \(x\)

OpenStudy (anonymous):

probably be easier to solve \[45x+25(8-x)=280\]for \(x\) because decimals are annoying

OpenStudy (anonymous):

@ktnguyen1 got the correct answer, but that was coincidence. just so happened the amounts were equal, which is what he/she assumed before writing the equation

OpenStudy (anonymous):

I see now...I ended up getting the 2.8, but was not sure if it was the correct answer. So this is only solving by one variable not two for x and y?

OpenStudy (anonymous):

Would solving for two variables be like my next problem: A car leaves St. Paul, Minnesota at 4 pm. A bus leaves the same place at 5 pm and travels in the same direction as the car. The car is traveling 12 mph faster than the bus. At 8 pm the vehicles are 100 miles apart. What is the speed of each car?

OpenStudy (anonymous):

i think we can do this

OpenStudy (anonymous):

lets call the rate of the car \(x\) so the rate of the bus is \(x-12\) because it is slower since distance is rate times time, we know the distance the car travelled is \(4x\) and the distance the bus travelled is \(3(x-12)\) the first distance is 100 miles more than the second distance, so ste \[4x=3(x-12)+100\] and solve for \(x\)

OpenStudy (anonymous):

this one is pretty easy to solve you get \[4x=3x-36+100\] \[4x=3x+64\] \[x=64\]

OpenStudy (anonymous):

Oh, ok I think I get it now. Thank you for your help!